OpenStudy (anonymous):

find all factors of x: x cube -12x+16=0 x cube-5x square+8=0

OpenStudy (anonymous):

\[x^{3}-12x+16=0 \]

OpenStudy (anonymous):

\[x^{3}-5x ^{2}+8=0\]

OpenStudy (anonymous):

@Hero

hero (hero):

You can find the first factor using trial and error. Afterwards, you can find the others by solving the remaining quadratic.

OpenStudy (anonymous):

ok like that helps..... sigh*

OpenStudy (anonymous):

hey hero

hero (hero):

\[(2)^3 - 12(2) + 16 \\= 8 - 24 + 16 \\= 8 + 16 - 24 \\= 24 - 24 \\= 0\] so so (x - 2) is a factor. Now, divide: \[\frac{x^3 - 12x + 16}{x - 2}\]. You will get a quadratic afterwards. Factor it to find the remaining factors.

OpenStudy (anonymous):

you factor the top one

hero (hero):

You have to divide first

hero (hero):

I explained what needed to be done.

OpenStudy (anonymous):

ok thanks for ya help even though i just cant get

hero (hero):

Polynomial long division

hero (hero):

@shanayduncan